# The Persistence of Contrarianism on Twitter: Mapping users’ sharing habits for the Ukraine war, COVID-19 vaccination, and the 2020 Midterm Elections

### v1

@@ Abstract
Empirical studies of online disinformation emphasize matters of public concern such as the COVID-19 pandemic, foreign election interference, and the Russo-Ukraine war, largely in studies that treat the topics separately. Comparatively fewer studies attempt to relate such disparate topics and address the extent to which they share behaviors. In this study, we compare three samples of Twitter data on COVID-19 vaccination, the Ukraine war and the 2020 midterm elections, to ascertain how distinct ideological stances of users across the three samples might be related. Our results indicate the emergence of a broad contrarian stance that is defined by its opposition to public health narratives/policies along with the Biden administration’s foreign policy stances. Sharing activity within the contrarian position falls on a spectrum with outright conspiratorial content on one end. We confirm the existence of ideologically coherent cross-subject stances among Twitter users, but in a manner not squarely aligned with right-left political orientations.
@@ 1 Introduction
In this study, we address these questions by comparing three samples of Twitter data on COVID-19 vaccination, the Ukraine war and the 2020 midterm elections. COVID-19 and Ukraine war are two of the largest global events of the last five years well represented on Twitter; the US 2020 midterm elections took place during the pandemic, and its inclusion alongside the other two helps situate them politically with respect to the US context. We adopt an approach employing Principal Components and cluster analysis to analyze the network of retweets, focusing in particular on a subset of users common to all three samples.
@@ 2 Data
This analysis employs two published Twitter datasets about COVID-19 vaccines [6] and the 2020 United States midterm elections [1] along with a third dataset on the Ukraine war that was collected using a similar method. For all three samples, data collection begins using a limited set of topic-relevant terms e.g., “Ukraine”; the initial collection is examined for additional relevant terms to expand the search query. The resulting datasets span January 4, 2021 to February 7, 2023 (for COVID-19), January 28, 2022 to December 2, 2022 (Ukraine), and September 10, 2022 to December 31, 2022 (midterm). Because these datasets contain varying amounts of spam and unrelated content, we further trimmed the dataset to retweets of persistently active users, identified by checking against a list of users active during October 2023. This results in retweets by 930,277 users in the COVID-19 data, 1,532,489 users in the Ukraine data, and 466,925 users in the midterm data.

### v2

# The Persistence of Contrarianism on Twitter: Mapping users’ sharing habits for the Ukraine war, COVID-19 vaccination, and the 2022 Midterm Elections

@@ Abstract
Empirical studies of online disinformation emphasize matters of public concern such as the COVID-19 pandemic, foreign election interference, and the Russo-Ukraine war, largely in studies that treat the topics separately. Comparatively fewer studies attempt to relate such disparate topics and address the extent to which they share behaviors. In this study, we compare three samples of Twitter data on COVID-19 vaccination, the Ukraine war and the 2022 midterm elections, to ascertain how distinct ideological stances of users across the three samples might be related. Our results indicate the emergence of a broad contrarian stance that is defined by its opposition to public health narratives/policies along with the Biden administration’s foreign policy stances. Sharing activity within the contrarian position falls on a spectrum with outright conspiratorial content on one end. We confirm the existence of ideologically coherent cross-subject stances among Twitter users, but in a manner not squarely aligned with right-left political orientations.
@@ 1 Introduction
In this study, we address these questions by comparing three samples of Twitter data on COVID-19 vaccination, the Ukraine war and the 2022 midterm elections. COVID-19 and Ukraine war are two of the largest global events of the last five years well represented on Twitter; the US 2022 midterm elections took place during the pandemic, and its inclusion alongside the other two helps situate them politically with respect to the US context. We adopt an approach employing Principal Components and cluster analysis to analyze the network of retweets, focusing in particular on a subset of users common to all three samples.
@@ 2 Data
This analysis employs two published Twitter datasets about COVID-19 vaccines [6] and the 2022 United States midterm elections [1] along with a third dataset on the Ukraine war that was collected using a similar method. For all three samples, data collection begins using a limited set of topic-relevant terms e.g., “Ukraine”; the initial collection is examined for additional relevant terms to expand the search query. The resulting datasets span January 4, 2021 to February 7, 2023 (for COVID-19), January 28, 2022 to December 2, 2022 (Ukraine), and September 10, 2022 to December 31, 2022 (midterm). Because these datasets contain varying amounts of spam and unrelated content, we further trimmed the dataset to retweets of persistently active users, identified by checking against a list of users active during October 2023. This results in retweets by 930,277 users in the COVID-19 data, 1,532,489 users in the Ukraine data, and 466,925 users in the midterm data.